OpenStudy (anonymous):

After 1 second, the ball is 121 feet in the air; after 2 seconds, the ball is 224 feet in the air. Find the height, in feet, of the ball after 3 seconds in the air. use the quadratic function to solve.

OpenStudy (anonymous):

I think you should indicate if there is initial velocity or not.

OpenStudy (anonymous):

The ball was launched from a sling shot. that's all i know.

OpenStudy (anonymous):

Sorry my stupid let me try to solve...

OpenStudy (anonymous):

it's okay. :b

OpenStudy (anonymous):

\[V0*t-1/2*g*t ^{2}\] V0*1-1/2*g*1=121 v0*2-1/2*g*4=224 After the solving equation , i found V0=130 feet/s , and g=18 feet/s^2 130*3-1/2*18*9=309 feet ,

OpenStudy (anonymous):

thanks! would you be interested in helping me with the last 3?

OpenStudy (anonymous):

If i can , but what you mean with last 3 ?

OpenStudy (anonymous):

i have 3 questions left.

OpenStudy (anonymous):

Ahh Okay (: i will try ... shoot!

OpenStudy (anonymous):

The company discovered that it costs $37 to produce 3 calculators, $213 to produce 7, $658 to produce 12. Using the quadratic function, find the cost of producing 8 calculators.

OpenStudy (anonymous):

Sorry i couldn't solve it. :/

OpenStudy (anonymous):

it's okay thanks though